OpenStudy (anonymous):

Beach Ball Problem -Francine has a beach ball. It is colored with six vertical stripes, in order: white, red, blue, yellow, orange and green. She spins the ball on the surface of the water in her swimming pool, and she notices that the colors whirl by her very fast. If the first color to go by is green, and the ball spins around so that 500 colors go by, what is the 500th color? Explain how you got your answer.

OpenStudy (anonymous):

6 colors white, red, blue, yellow, orange green which way she spin the ball? cuzz if she spin -> then next color is white but if she spin <- next color is orange

OpenStudy (anonymous):

->

OpenStudy (anonymous):

ok thanks this helps

OpenStudy (anonymous):

6 colors white, red, blue, yellow, orange green spins just to keep track. 2 3 4 5 6 1 1 8 9 10 11 12 7 2 14 15 16 17 18 13 3 green1= 1 green2= green1+6 green3= green2+6

OpenStudy (anonymous):

1 =1 0*6 1 spin 1 +6 =7 1*6 2 spin (1+6)+6 = 13 2*6 3 spin 1+6+6+6= 19 3*6 4 spin

OpenStudy (anonymous):

so at ith spin green will be 1+6*(i-1) lets tri to find out roughly when green will be 500 500+a=1+6*(i-1) the a will be cuzz green wont be exactly the 500th color

OpenStudy (anonymous):

500+a=1+6*(i-1) or if we choose a = 0 then 500=1+6*(i-1) sfi i = 84.1 so lets choose 85th spin 1+6*(85-1) = 505 will be green how about 84th spin 1+6*(84-1)=499. :D which means the next color is the 500th which is WHITE

OpenStudy (anonymous):

oh my gosh dude thank you ive been stumped on this question!

OpenStudy (anonymous):

:D just do it like me. ANd plot and draw and make tables :D

OpenStudy (anonymous):

not to confuse you but i think this might work or not. 1+6*(i-1) ^ that is first and green but white is 2 and red is 3 and so on so lets say a+6*(i-1) i think that is true. a+6*(i-1)=500 you'd have to find an a and i. An a from 1 to 6 But it looks like it might take longer lol. nvrmnd